which choice is equivalent to the fraction below? Hint: rationalize the denominator and simplify.
2-2√2/2+2√2
a. 1-√2
b. -3+2√2
c. 2+√2
d. 2-√2

Answer by edjones(8007)

2-2√2/2+2√2
=2(1-sqrt(2))/2(1+sqrt(2))
=(1-sqrt(2))/(1+sqrt(2))
=
=(1-sqrt(2))^2/(1-2)
=(1-2sqrt(2)+2)/-1
=-3+2sqrt(2)
.
Ed